OpenStudy (naga447):

Simplifying help!!!! square root of 2 multiplied by the cube root of 2

OpenStudy (igreen):

What do you need it simplified to? @naga447

OpenStudy (igreen):

\(\sqrt{2} * 3\sqrt2\) = 1.78

OpenStudy (anonymous):

\(\sqrt2\sqrt[3]2 = 2^{\frac{1}{2}}\cdot2^{\frac{1}{3}}\) I think it is clear what we need it simplified to @iGreen

OpenStudy (igreen):

Oh, if you're doing that.. you would end up with \(2^\frac{5}{6}\)

OpenStudy (anonymous):

yeah, look simplified than \(\sqrt2\sqrt[3]2\)

OpenStudy (naga447):

Thank you both so much!~ @micahwood50 @iGreen